f(x)= (x-4)squared -1
Answer by 2897696(96)

a negative exponent changes the thing in the () to a fraction:
ex
: a^-1 ==> 1/(a^1)
: (1/a)^-1 ==> a^1
f(x) =(x-4)^-1 becomes: f(x)= 1/(x-4)